If `a` and `b` are natural numbers such that (1/a)^(1/b)=0.3 or [{one divided by a to the power of 1/b}]Find ab.

Grade:12

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
Given a,b are natural numbers
(1/a)(1/b)=0.3
taking log on both sides we get
1/b(log 1/a)=log 0.3
on solving we get a=100.523b
if b=1 a=4.3 which is not a natural number
if b=2,a=11 which is a matural number
So the value of ab=11*2=22
Thanks and Regards
Shaik Aasif
askIITians faculty